# https://vl2g.github.io/projects/IndicPhotoOCR/

### v1

@@ 2.4 Dataset Analysis
In this section, we present a comprehensive analysis of the Bharat Scene Text Dataset (BSTD) to evaluate its quality and the challenges it poses. Toward this end, we first report the language-wise word count in Table 3. We observe that it has a good coverage of all the major Indian languages with minimum 2,542 word images for Tamil. This makes the dataset linguistically rich. We further analyze the average word length for each language in our dataset in Fig. 5). The plot reveals that Malayalam, Tamil, and Telugu contain longer words on average, whereas English and Hindi have shorter word lengths. This observation, though potentially intuitive, holds practical implications for the development of specialized language models in the future.
Moving further, we studied the distribution of the bounding box area in Fig. 6. It indicates that most text regions (59.4%) are small (1001–10,000 pixels), while only 2.6% of text instances fall into the largest category (above 100,000 pixels). These findings highlight the diverse textual characteristics within BSTD and it contains a mix of small and large text regions within images making it challenging for detection. Furthermore, the high percentage of small text regions suggests that many images contain incidental text rather than just focused text, which adds to the dataset’s difficulty implying the real word scenario.
Figure 7: Word cloud for recognition annotations in our BSTD dataset: Highlighting the diversity of vocabulary across 12 languages. The tiles are arranged row-wise from left to right as follows: the first row contains Hindi, Assamese, and Gujarati; the second row shows English, Malayalam, and Kannada; the third row corresponds to Odia, Panjabi, and Tamil; and the fourth row includes Telugu, Marathi, and Bengali.
Figure 8: A selection of challenging cropped words from our dataset. Few key challenges include eroded text (c, d, k), handwritten text (f), occluded text (e, h, l), ambiguous text (a, b, i), text in different orientations (g, j).
Finally, to illustrate diversity in the dataset, we show word clouds for all 12 languages in our data set in Fig. 7. As can be seen, the most frequent words (represented by the largest font size) across all languages point to text commonly found on public signs, government buildings, and commercial establishments. In addition, the vocabulary is dominated by nouns that are staples of scene text recognition tasks. These are words that one would expect to see “in the wild”. Some of the cropped word samples from BSTD samples are shown in Fig. 8.
@@
In short, the analysis of the Bharat Scene Text Dataset suggests that it has diverse linguistic and geographical coverage, with a clear focus on text from public, commercial and transport-related signage, and has a mix of focused and incidental text, making it a valuable resource for developing robust real-world text recognition models for Indian languages. Further, while widely spoken languages such as Hindi, English, Bengali, and Marathi are well represented in the dataset, low-resource languages including Assamese, Gujarati, Kannada, Malayalam, Odia, Punjabi, Tamil, and Telugu also have a substantial presence, with at least 2,542 word images for Tamil (see Table 3). This shows that the dataset is not only strong in high-resource languages but also sufficiently representative of low-resource languages, making it suitable for comprehensive multilingual text recognition research.
@@
Scene text recognition in a multilingual environment typically involves three key modules: (i) Scene Text Detection, (ii) Script Identification, and (iii) Cropped Word Recognition. In the absence of a dedicated architecture for Indian language scene text recognition in the literature, we develop a strong baseline by adapting or fine-tuning (as needed) state-of-the-art architectures across various modules. We describe each of these modules below and collectively refer to this pipeline as [Uncaptioned image].
@@ 3.1 Scene Text Detection
Our systematic analysis of modern scene text detectors revealed that these models are largely language-agnostic. Hence, we directly incorporate them into our pipeline without language-specific modifications. In particular, we adopt TextBPN++ [24] for scene text detection, as it demonstrated superior performance compared to other detectors in our evaluations. Detailed results for this stage are presented in the experimental section 4.1.
@@ 3.2 Script Identification
We fine-tuned the model, including both the transformer backbone and the classification head, using the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a batch size of 16 using BSTD-train set. This ViT-based approach consistently outperformed other configurations, demonstrating strong generalization in both script classification settings of the 3-way and 12-way scripts.
Language Train Size Test Size Vocabulary Size #Fonts
@@
Table 4: Statistics for Synthetic Dataset generated using [36] for 11 Indian languages.
@@ 3.3 Cropped Word Recognition
Once text instances are detected and their scripts identified, the next step is text recognition. Unlike detection, this task relies heavily on language-specific modeling and requires large-scale training data. Off-the-shelf models for English text recognition cannot be directly applied as they are not designed for this task. Instead, we fine-tune a successful scene text recognition architecture (namely PARSeq [22]) for the 11 Indian languages in our study. A key challenge in this setting is the scarcity of real-world scene text data for low-resource languages, in contrast to English, where recognizers are trained on millions of samples. To address this, we generated synthetic word images to augment the training data and bridge this gap as described in the following section.
@@ 3.3.1 Synthetic Data Generation
To train models for each of the 11 languages, we required a large-scale dataset of scene text images. However, such datasets for Indian languages were not readily available. This lack of data led us to adopt an existing method for generating synthetic scene text images. We used SynthText [36] to generate large-scale data, with statistics summarized in Table 4. SynthText [36] provided 8000 background images, which we leveraged as the background image for our synthetic data.
To generate scene text in different Indian languages, we needed appropriate vocabulary and fonts for each language. We sourced vocabulary from the AI4Bharat-IndicNLP Dataset [^7] https://github.com/AI4Bharat/indicnlp_corpus introduced in the IndicNLPSuite [29]. SynthText also accounts for local 3D geometry, enabling varied orientation of text placement within the scene. Although the generated images lacked complete realism, the scale of data generation offered significant value for our training needs.
@@ 4 Experimental Analysis
In this section, we provide a detailed evaluation of the proposed baseline, namely [Uncaptioned image], and its individual modules on the newly introduced Bharat Scene Text Dataset. We compare it with several open-source OCR systems, including Tesseract [37], PaddleOCR [38], EasyOCR [39], and SuryaOCR [40]. For end-to-end evaluation, we also compare it with commercial closed-source systems such as Google OCR [41] and GPT-4 [42]. In addition, we include appropriate competitive approaches to evaluate individual modules. Our results reveal both the strengths and limitations of existing systems in handling diverse multilingual Indian scene text. They also demonstrate the effectiveness of [Uncaptioned image] in a challenging setting. Please note that all the results are reported on the Bharat Scene Text Dataset (test set).
@@ 4.1 Results on Scene-text detection
Our proposed baseline, [Uncaptioned image], employs TextBPN++ [24] for scene text detection. We compare its performance with several widely used scene text detection methods, including EAST [14], CRAFT [16], DBNet [19], and Hi-SAM [43]. The comparison results are summarized in Table 5. We observe that TextBPN++ significantly outperforms the other detectors, achieving higher F1-score and recall. Although Hi-SAM is also equally precise, it misses out many in detecting many Indian language words. Please note that we use TedEval [35] evaluation for reporting these numbers.
Method P R F1
EAST [14] 0.55 0.10 0.17
CRAFT [16] 0.51 0.22 0.19
DBNet [19] 0.71 0.50 0.59
Hi-SAM [43] 0.75 0.19 0.46
TextBPN++ [24] 0.75 0.78 0.77
4.2 Results on Script Identification
We use the following baselines: (i) AlexNet [44]: a popular CNN-based Imagenet-pretrained image classification model finetuned on trainset of our script identification data, (ii) CRNN [45]: a CNN-based scene text recognition model with the ability to encode sequence information, adapted for script identification, followed by state-of-the-art transformer architectures for image classification; (iii) CLIP [46] and (iv) ViT [47], respectively. All the results reported in Table 6 correspond to the models trained on real images.
@@ 4.2 Results on Script Identification
Language AlexNet [44] CRNN [12] CLIP [46] Proposed
Assamese 53.1 72 75.6 92.2
Bengali 54 75.7 77.5 93.3
@@
Table 6: Script Identification Accuracy in 3-way classification setting on Hindi vs English vs third regional Indian language shown in the language column.
Figure 9: Normalized confusion matrix illustrating the performance of script identification model across 12 Indian. (Best viewed in color).
@@ 4.3 Results on Cropped Word Recognition
For the cropped-word recognition baselines, there are no models specifically designed for Indian languages. As a result, we must rely on English scene-text recognition approaches such as [25, 48, 22, 45]. However, adopting all these methods still demands substantial computational resources, especially when running recognition models across 12 languages. Therefore, we only choose PARSeq [22], one of the recent and effective recognition models, as our baseline. In addition, we compare the PARSeq-based baseline with the already available, CRNN [12] by AI4Bharat [49] and off-the-shelf open-source models, namely Tesseract [37], PaddleOCR [38], and EasyOCR [39].
@@
ParSeq is trained on a combination of synthetic data and the BSTD dataset. Our baseline for cropped word recognition is based on the existing text recognition framework PARSeq [22]. We report results under two settings: (i) training PARSeq from scratch using synthetic data, and (ii) fine-tuning it on the BSTD training set. The results are presented in Table 7. Fine-tuning on real data leads to substantial improvements in recognition accuracy. This highlights the critical role of real-world data in scene text recognition. Also, as can be seen, Word Recognition Rates for cropped English scene text are relatively high (around 92%). However, performance drops noticeably for other languages: Marathi (86%), Bengali (82%), Tamil (80%), Aasamease (79%) Punjabi (75%), Odia (72%), and Hindi (71%) form the next tier of top-performing languages. The gap widens further for low-resource languages such as Telugu and Malayalam, whose recognition accuracies remain in the 50% range. These results clearly indicate substantial room for improving recognition performance across Indian languages.
In the next section we will be evaluating recognition performance in even more challenging end-to-end setting.
PARSeq [22] Synth 0.47 0.44 0.92 0.32 0.43 0.47 0.52 0.44 0.42 0.4 0.37 0.44 0.47
BSTD Train 0.79 0.82 0.92 0.61 0.71 0.6 0.58 0.86 0.72 0.75 0.8 0.56 0.73
CRNN [12] AI4Bharat NA NA 0.27 NA 0.13 NA 0.06 NA NA 0.27 0.16 0.05 0.16
Off-the-shelf open source models
Tesseract [37] - 0.15 0.18 0.33 0.12 0.2 0.12 0.03 0.21 0.11 0.19 0.13 0.09 0.15
PaddleOCR [38] - NA NA 0.61 NA 0.31 0.03 NA 0.36 NA NA 0.28 0.15 0.29
EasyOCR [39] - 0.12 0.24 0.29 NA 0.18 0.06 NA 0.25 NA NA NA 0.11 0.18
Table 7: Baseline Results on Cropped Word Scene Text Recognition for 12 languages on BSTD. Please note that PARSeq in used in our proposed baseline – [Uncaptioned image].
WRR on BSTD-Test
Method /b Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
ChatGPT [42] 0.08 0.36 0.09 0.2 0.03 0.05 0.04 0.12 0.21 0.1 0.13
Google OCR [41] 0.43 0.37 0.32 0.55 0.24 0.37 0.43 0.46 0.52 0.42 0.41
Tesseract [37] b 0.04 0.02 0.02 0.03 0 0 0.02 0.02 0 0.02 0.02
SuryaOCR [40] b 0.18 0.15 0.09 0.19 0.07 0.12 0.15 0.21 0.24 0.05 0.14
[Uncaptioned image] b 0.45 0.37 0.3 0.39 0.29 0.31 0.38 0.41 0.42 0.31 0.36
[Uncaptioned image] (+oracle TD) b 0.75 0.74 0.54 0.63 0.49 0.54 0.68 0.7 0.69 0.49 0.62
[Uncaptioned image] (+oracle TD+SI) b 0.8 0.92 0.61 0.74 0.6 0.58 0.72 0.75 0.8 0.56 0.71
CRR on BSTD-Test
Method /b Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
ChatGPT [42] 0.13 0.46 0.15 0.28 0.1 0.14 0.11 0.2 0.31 0.21 0.21
Google OCR [41] 0.57 0.42 0.36 0.66 0.45 0.55 0.57 0.59 0.69 0.6 0.55
Tesseract [37] b 0.12 0.07 0.06 0.09 0.05 0.09 0.09 0.09 0.04 0.06 0.08
SuryaOCR [40] b 0.29 0.24 0.15 0.33 0.19 0.26 0.29 0.31 0.45 0.26 0.28
[Uncaptioned image] b 0.59 0.47 0.43 0.54 0.48 0.55 0.59 0.53 0.62 0.57 0.54
[Uncaptioned image] (+oracle TD) b 0.87 0.78 0.64 0.74 0.68 0.8 0.84 0.84 0.78 0.73 0.77
[Uncaptioned image] (+oracle TD+SI) b 0.93 0.97 0.75 0.89 0.86 0.86 0.89 0.9 0.94 0.83 0.88
@@
Table 8: We conduct a comparative evaluation of commercial systems, including Google OCR and GPT-4, alongside our proposed baseline namely [Uncaptioned image] for end-to-end scene text recognition on the BSTD dataset. The evaluation measures script-wise performance using Word Recognition Rate (WRR) and Character Recognition Rate (CRR). Additionally, we present baseline results under two scenarios: (i) when text detection is provided as ground truth (denoted as oracle TD) and (ii) when both text detection and script identification are given as ground truth (denoted as TD+SI). “ /b” indicates whether the tool is open-source ( b) or closed-source (). Bold and underline number shows best performing results in non-oracle setting for closed and open-source models respectively. [Uncaptioned image] is the best open-source model for all the Indian languages. The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
Method /b Metric Languages Avg.
@@
[Uncaptioned image](+oracle TD) b P 0.72 0.83 0.65 0.66 0.56 0.57 0.65 0.69 0.78 0.54 0.66
@@
[Uncaptioned image](+oracle TD+SI) b P 0.75 0.89 0.71 0.71 0.57 0.6 0.66 0.72 0.8 0.55 0.70
@@
Table 9: Performance comparison across 10 Indic scripts on the BSTD dataset for end-to-end scene text recognition. We report Precision (P), Recall (R), and F1-score (F) for each script and the average performance. The proposed baseline model, [Uncaptioned image], significantly outperforms available open-source models and achieves competitive performance compared to commercial models. The inclusion of oracle Text Detection (TD) and Script Identification (SI) shows the potential for further improvement. The best scores for closed-source and open-source are shown in bold and underlined, respectively. The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.

### v2

@@ 2.4 Dataset Analysis
In this section, we present a comprehensive analysis of the Bharat Scene Text Dataset (BSTD) to evaluate its quality and the challenges it poses. Toward this end, we first report the language-wise word count in Table 3. We observe that it has a good coverage of all the major Indian languages with minimum 2,542 word images for Tamil. Further, we performed an image-level analysis and observed that, out of the 6,582 total scene images in the dataset, only 53 contain exclusively English text, 1,958 contain only non-English text (i.e., Indian languages), and 4,518 contain English co-occurring with one or more Indian languages. This makes the dataset linguistically rich (refer to Table 4). We further analyze the average word length for each language in our dataset in Fig. 5. The plot reveals that Malayalam, Tamil, and Telugu contain longer words on average, whereas English and Hindi have shorter word lengths. This observation, though potentially intuitive, holds practical implications for the development of specialized language models in the future.
Moving further, we studied the distribution of the bounding box area in Fig. 6. It indicates that bounding box sizes range from very small text region, e.g., minimum area = 32 pixels to very-large text region, e.g., maximum area $\approx$ $2.49\times 10^{6}$ pixels. The distribution closely approximates a log-normal curve that peaks near the median area of 2,889 pixels. These findings highlight that the BSTD contains a mix of small and large text regions within images, making it challenging for detection. Furthermore, the high proportion of small text regions indicates the prevalence of incidental text, as opposed to only focused text, thereby making the dataset more challenging and representative of real-world scenarios.
An analysis of the lighting and environmental conditions of the scene images revealed that approximately 89% of the images were captured during daylight, whereas the remaining 11% were taken at night. Further, the environment annotations indicate that 91% of the images correspond to outdoor scenes, while only 9% were captured indoors.
@@
In short, the analysis of the Bharat Scene Text Dataset suggests that it has diverse linguistic and geographical coverage, with a clear focus on text from public, commercial and transport-related signage, and has a mix of focused and incidental text, making it a valuable resource for developing robust real-world text recognition models for Indian languages. Further, while widely spoken languages such as Hindi, English, Bengali, and Marathi are well represented in the dataset, low-resource languages including Assamese, Gujarati, Kannada, Malayalam, Odia, Punjabi, Tamil, and Telugu also have a substantial presence, with at least 2,542 word images for Tamil (see Table 3). This shows that the dataset is not only strong in high-resource languages, but also sufficiently representative of low-resource languages, making it suitable for comprehensive multilingual text recognition research.
@@
Scene text recognition in a multilingual environment typically involves three key modules: (i) Scene Text Detection, (ii) Script Identification, and (iii) Cropped Word Recognition. In the absence of a dedicated architecture for Indian language scene text recognition in the literature, we develop a strong baseline by adapting or fine-tuning (as needed) state-of-the-art architectures across these modules. We describe each of these modules below and collectively refer to this pipeline as [Uncaptioned image].
@@ 3.1 Scene Text Detection
Our systematic analysis of modern scene text detectors revealed that these models are largely language-agnostic. Hence, we directly incorporate them into our pipeline without language-specific modifications. In particular, we adopt TextBPN++ [textbpn++] for scene text detection, as it demonstrated superior performance compared to other detectors in our evaluations. Detailed results for this stage are presented in the experimental section 4.1.
@@ 3.2 Script Identification
We fine-tuned the model, including both the transformer backbone and the classification head, using the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a batch size of 16 using BSTD-train set. This ViT-based approach consistently outperformed other configurations, demonstrating strong generalization in both 3-way and 12-way script classification settings.
Table 5: Statistics for Synthetic Dataset generated using [synthtext] for 11 Indian languages.
Language Train Size Test Size Vocab Size #Fonts
@@ 3.3 Cropped Word Recognition
Once text instances are detected and their scripts identified, the next step is text recognition. In contrast to detection, this task depends strongly on language-specific modeling and demands large-scale training data. Existing off-the-shelf English scene text recognition models are not directly applicable, as they are not designed for Indian scripts. To address this, we fine-tune a successful scene text recognition architecture, PARSeq [parseq], for the 11 Indian languages included in our study. A key challenge in this setting is the scarcity of real-world scene text data for low-resource languages, in contrast to English, where recognizers are trained on millions of samples. To address this, we generated synthetic word images to augment the training data and bridge this gap as described in the following section.
@@ 3.3.1 Synthetic Data Generation
To train models for each of the 11 languages, we required a large-scale dataset of scene text images. However, such datasets for Indian languages were not readily available. This lack of data led us to adopt an existing method for generating synthetic scene text images. We used SynthText [synthtext] to generate large-scale data, with statistics summarized in Table 5. SynthText [synthtext] provided 8000 background images, which we leveraged as the background image for our synthetic data.
To generate scene text in different Indian languages, we needed appropriate vocabulary and fonts for each language. We sourced vocabulary from the AI4Bharat-IndicNLP Dataset [^7] https://github.com/AI4Bharat/indicnlp_corpus introduced in the IndicNLPSuite [kakwani-etal-2020-indicnlpsuite]. SynthText also accounts for local 3D geometry, enabling varied orientation of text placement within the scene. Although the generated images lacked complete realism, the scale of data generation offered significant value for our training needs.
@@ 4 Experimental Analysis
In this section, we provide a detailed evaluation of the proposed baseline, namely [Uncaptioned image], and its individual modules on the newly introduced Bharat Scene Text Dataset. We compare it with several open-source OCR systems, including Tesseract [tesseract], PaddleOCR [cui2025paddleocr30technicalreport], EasyOCR [jaidedai_easyocr_2020], and SuryaOCR [paruchuri2025surya]. For end-to-end evaluation, we also compare it with commercial closed-source systems such as Google OCR [googlecloud_visionocr] and GPT-4 [gpt_4v]. In addition, we include appropriate competitive approaches to evaluate individual modules. Our results reveal both the strengths and limitations of existing systems in handling diverse multilingual Indian scene text. They also demonstrate the effectiveness of [Uncaptioned image] in a challenging setting. Please note that all the quantitative results are reported on the Bharat Scene Text Dataset (test set).
Method P R F1
EAST [east] 0.55 0.10 0.17
CRAFT [craft] 0.51 0.22 0.19
DBNet [dbnet] 0.71 0.50 0.59
Hi-SAM [Hi-SAM] 0.75 0.19 0.46
TextBPN++ [textbpn++] 0.75 0.78 0.77
@@ 4.1 Results on Scene-text detection
Our proposed baseline, [Uncaptioned image], employs TextBPN++ [textbpn++] for scene text detection. We compare its performance with several widely used scene text detection methods, including EAST [east], CRAFT [craft], DBNet [dbnet], and Hi-SAM [Hi-SAM]. The comparison results are summarized in Table 6. We observe that TextBPN++ significantly outperforms the other detectors, achieving higher F1-score and recall. Although Hi-SAM is also equally precise, it misses out detecting many Indian language words. Please note that we use TedEval [tedeval] evaluation in our experiments.
Table 7: Script identification accuracy (%) in the 3-way classification setting: Hindi vs English vs one regional Indian language shown in the language column. Each row corresponds to a different regional Indian language, while Hindi and English remain fixed across all rows.
@@
Language AlexNet CRNN CLIP Proposed
Assamese 53.1 72.0 75.6 92.2
Bengali 54.0 75.7 77.5 93.3
@@
4.2 Results on Script Identification
We use the following baselines: (i) AlexNet [alexnet]: finetuned on trainset of our script identification data, (ii) CRNN [CRNN2016]: a CNN-based scene text recognition model with the ability to encode sequence information, adapted for script identification, followed by state-of-the-art transformer architectures for image classification; (iii) CLIP [clip] and (iv) ViT [vit], respectively.
Figure 7: Normalized confusion matrix illustrating the performance of script identification in 12-way classification setting. (Best viewed in color).
@@ 4.2 Results on Script Identification
Table 8: Baseline Results on Cropped Word Scene Text Recognition for 12 languages on BSTD. Please note that PARSeq is used in our proposed baseline – [Uncaptioned image].
PARSeq [parseq] Synth 0.47 0.44 0.92 0.32 0.43 0.47 0.52 0.44 0.42 0.4 0.37 0.44 0.47
BSTD Train 0.79 0.82 0.92 0.61 0.71 0.6 0.58 0.86 0.72 0.75 0.8 0.56 0.73
CRNN [CRNN2016] AI4Bharat NA NA 0.27 NA 0.13 NA 0.06 NA NA 0.27 0.16 0.05 0.16
Off-the-shelf open source models
Tesseract [tesseract] - 0.15 0.18 0.33 0.12 0.2 0.12 0.03 0.21 0.11 0.19 0.13 0.09 0.15
PaddleOCR [cui2025paddleocr30technicalreport] - NA NA 0.61 NA 0.31 0.03 NA 0.36 NA NA 0.28 0.15 0.29
EasyOCR [jaidedai_easyocr_2020] - 0.12 0.24 0.29 NA 0.18 0.06 NA 0.25 NA NA NA 0.11 0.18
Figure 8: A selection of challenging cropped words from our dataset along with ground truth and recognized results by various methods including Refer to caption. Key challenges include out-of-focus text (a, b, i), varied fonts (c), eroded text (d, k), handwritten text (f), occluded text (e, h, l), and skewed text (g, j). Cells marked with “-” indicate an empty recognition, while “NA” denotes that specific language support was unavailable for that model. Here, OCR output text in red color shows that the word is incorrectly recognized. (Best viewed in color).
Table 9: Comparative evaluation for end-to-end scene text recognition on BSTD using WRR and CRR. We present baseline results under two scenarios: (i) when text detection is provided as ground truth (denoted as oracle TD) and (ii) when both text detection and script identification are given as ground truth (denoted as TD+SI). “ / ” indicates whether the tool is open-source ( ) or closed-source ( ). Bold and underlined number shows best performing results in non-oracle setting for closed and open-source models respectively. The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
WRR on BSTD-Test
Method / Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
ChatGPT [gpt_4v] 0.08 0.36 0.09 0.2 0.03 0.05 0.04 0.12 0.21 0.1 0.13
Google OCR [googlecloud_visionocr] 0.43 0.37 0.32 0.55 0.24 0.37 0.43 0.46 0.52 0.42 0.41
Tesseract [tesseract] 0.04 0.02 0.02 0.03 0 0 0.02 0.02 0 0.02 0.02
SuryaOCR [paruchuri2025surya] 0.18 0.15 0.09 0.19 0.07 0.12 0.15 0.21 0.24 0.05 0.14
[Uncaptioned image] 0.45 0.37 0.3 0.39 0.29 0.31 0.38 0.41 0.42 0.31 0.36
[Uncaptioned image] (+oracle TD) 0.75 0.74 0.54 0.63 0.49 0.54 0.68 0.7 0.69 0.49 0.62
[Uncaptioned image] (+oracle TD+SI) 0.8 0.92 0.61 0.74 0.6 0.58 0.72 0.75 0.8 0.56 0.71
CRR on BSTD-Test
Method / Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
ChatGPT [gpt_4v] 0.13 0.46 0.15 0.28 0.1 0.14 0.11 0.2 0.31 0.21 0.21
Google OCR [googlecloud_visionocr] 0.57 0.42 0.36 0.66 0.45 0.55 0.57 0.59 0.69 0.6 0.55
Tesseract [tesseract] 0.12 0.07 0.06 0.09 0.05 0.09 0.09 0.09 0.04 0.06 0.08
SuryaOCR [paruchuri2025surya] 0.29 0.24 0.15 0.33 0.19 0.26 0.29 0.31 0.45 0.26 0.28
[Uncaptioned image] 0.59 0.47 0.43 0.54 0.48 0.55 0.59 0.53 0.62 0.57 0.54
[Uncaptioned image] (+oracle TD) 0.87 0.78 0.64 0.74 0.68 0.8 0.84 0.84 0.78 0.73 0.77
[Uncaptioned image] (+oracle TD+SI) 0.93 0.97 0.75 0.89 0.86 0.86 0.89 0.9 0.94 0.83 0.88
@@ 4.3 Results on Cropped Word Recognition
For the cropped-word recognition baselines, there are no models specifically designed for Indian languages. Therefore, we could adopt English scene-text recognition approaches such as [clip4str, clipter, parseq, CRNN2016]. However, adopting all these methods still demands substantial computational resources, especially when running recognition models across 12 languages. Therefore, we only choose PARSeq [parseq], one of the recent and effective recognition models, as our baseline. In addition, we also use available, CRNN [CRNN2016] by AI4Bharat [gokulkarthik_indianstr_2023] and off-the-shelf open-source models, namely Tesseract [tesseract],PaddleOCR [cui2025paddleocr30technicalreport], and EasyOCR [jaidedai_easyocr_2020].
@@
PARSeq is trained on a combination of synthetic data and the BSTD dataset. Our baseline for cropped word recognition is based on the existing text recognition framework PARSeq [parseq]. We report results under two settings: (i) training PARSeq from scratch using synthetic data, and (ii) fine-tuning it on the BSTD training set. The results are presented in Table 8. Fine-tuning on real data leads to substantial improvements in recognition accuracy. This highlights the critical role of real-world data in scene text recognition. Also, as can be seen, Word Recognition Rates for cropped English scene text are relatively high (around 92%) [^8] Unless otherwise specified, the English recognition model is trained on both uppercase and lowercase characters; however, the decoder normalizes predictions to lowercase, and evaluation is performed in a case-insensitive manner. We additionally conducted an experiment without this relaxation, i.e., using case-sensitive decoding and evaluation. Under this stricter setting, the model achieves a WRR of 84%, compared to 92% under case-insensitive decoding.. However, performance drops noticeably for other languages: Marathi (86%), Bengali (82%), Tamil (80%), Aasamease (79%) Punjabi (75%), Odia (72%), and Hindi (71%) form the next tier of top-performing languages. The gap widens further for low-resource languages such as Telugu and Malayalam, whose recognition accuracies remain in the 50% range. These results clearly indicate substantial room for improving recognition performance across Indian languages.
Futhermore, we show a selection of visual results for several challenging cropped word images from BSTD, comparing predictions from our model against the baseline in Fig. 8.
@@
Table 10: Comparative evaluation for end-to-end scene text recognition on BSTD using Precision (P), Recall (R) and F-score (F). . The proposed baseline model, [Uncaptioned image], significantly outperforms available open-source models and achieves competitive performance compared to commercial models. The inclusion of oracle Text Detection (TD) and Script Identification (SI) shows the potential for further improvement. The best scores for closed-source and open-source are shown in bold and underlined, respectively. The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
Method / Metric Languages Avg.
@@
[Uncaptioned image](+oracle TD) P 0.72 0.83 0.65 0.66 0.56 0.57 0.65 0.69 0.78 0.54 0.66
@@
[Uncaptioned image](+oracle TD+SI) P 0.75 0.89 0.71 0.71 0.57 0.6 0.66 0.72 0.8 0.55 0.70